OpenStudy (jmark):

Hi can you please help me to solve this prob? Find the number divisors of 240

OpenStudy (alexandervonhumboldt2):

the first divisor is 2: 240/2=120 120/2=60 60/2=30 30/2=15 15/3=5 5/5=1 each time you divide new number by the it's smallest divisor: thus your divisors are 2, 4, 8, 16, 6, 10, 48, 30, 15, 60, 5, 120, 240, 1, 3, 80 , 24, 12, 20, 40, that's all if i didn't forgot anything prime factorization is 2*2*2*2*3*5

OpenStudy (alexandervonhumboldt2):

to find divisors of 240, first write it's prime factorization, then calculate all of the combinations between the prime factors

OpenStudy (alexandervonhumboldt2):

@JMark

OpenStudy (mom.):

what kind of divisors? 2.4 is also a divisor :)

OpenStudy (mathmath333):

is the answer 20

ganeshie8 (ganeshie8):

\(240 = 2^4*3*5\) So, any divisor of \(240\) will be of form \(2^a*3^b*5^c\), where \(0\le a\le 4\) and \(~0\le b,c\le 1\)

OpenStudy (alexandervonhumboldt2):

yeah gane is right

OpenStudy (mathmath333):

mom=Mars Orbiter Mission

OpenStudy (mom.):

mom=mother of mankind (;

ganeshie8 (ganeshie8):

Notice that we can choose the exponent \(a\) in \(5\) ways, \(b\) in \(2\) ways, \(c\) in \(2\) ways therefore total number of positive divisors = \(5*2*2=20\)

OpenStudy (mathmath333):

i learnt this stuff 1 year ago here now i forgot it

OpenStudy (mom.):

same

OpenStudy (alexandervonhumboldt2):

looks like Mister.Mark is not in a hurry to see our explanations @JMark

OpenStudy (mathmath333):

i just remembered the pattern (a+1)(b+1)(c+1)....=no of divisors

OpenStudy (mom.):

looks like he did it himself (;